OpenStudy (anonymous):

Finding eigenvectors: Can some just show me the steps to find the eigenvector of the following matrix, with (lambda)=2. Matrix is given below.

OpenStudy (anonymous):

\[\left[\begin{matrix}-1 & 2 & -1\\ 3 & 0 & 1\\ -3 & -2 & -3\end{matrix}\right]\]

OpenStudy (anonymous):

I hated linear algebra and don't remember anything, but I will suggest Paul's online notes for these (just Google is). Also, if you can get your hands on a pdf copy of 'Elementary Linear Algebra' by Larson and Falvo.

OpenStudy (zzr0ck3r):

first do A - bI where b is some constant variable and I is the identity matrix

OpenStudy (zzr0ck3r):

so [-1-b,2,-1;3,-b,1;-3,-2,-3-b]

OpenStudy (zzr0ck3r):

understand? b is your lamda

OpenStudy (anonymous):

yes, but I thought you used A-bI=0, in order to find all the lambdas... and then use Ax=bx, in order to find the eigenvectors.

OpenStudy (zzr0ck3r):

o sorry you labda is given

OpenStudy (zzr0ck3r):

you have your lambda its 2

OpenStudy (anonymous):

and I have all three eigenvalues, but for some reason when i am calculating for the eigenvector, i am not getting the right answer

OpenStudy (zzr0ck3r):

so solve for the null space of that equation

OpenStudy (zzr0ck3r):

one sec

OpenStudy (anonymous):

and if i can see how to do it just for lambda = 2 , i can figure out the other two eigenvectors.

OpenStudy (turingtest):

if I remember this right, you just plug in each lambda and solve (A-b_1*I)=0 (A-b_2*I)=0

OpenStudy (anonymous):

ahhh, got it, okay... i see where i went wrong, thank you for the help

OpenStudy (zzr0ck3r):

np